Prove that the set of complex numbers {1,−1,i,−i} under

multiplication operation is a cyclic group.Find the generators of

cycle


1
Expert's answer
2022-04-18T00:08:55-0400

The group that contains a finite number of elements is cyclic, in case there is an element "g" of this group such that all elements have the form: "g^j", where "j\in{\mathbb{N}}". I.e., all elements can be obtained from "g". In case of set "\{1,-1,i,-i\}" we have: "i^2=-1," "i^3=-i", "i^4=1". Thus, all elements can be recovered from "i". The group is cyclic. "i" is a generator of the group. Consider other elements of the group: "(-i)^2=-1", "(-i)^3=i", "(-i)^4=1". . "(-1)^2=1", "1^2=1."

Thus, the answer is: the group is cyclic, the generators are "(-i)" and "i". "(-1)" generates a subgroup: "\{-1,1\}". "1" generates only itself.